# Understanding Benefits and Pitfalls of Current Methods for the Segmentation of Undersampled MRI Data

### v1

@@ 2 Related Work
Predicting high-quality segmentations has been the objective in the K2S challenge (Tolpadi et al., 2023), where several methods for segmenting undersampled MRI data have been submitted. While reconstruction fidelity was also analyzed for the methods that did provide reconstructions, it was not used as an evaluation criterion. A drawback of the challenge is that only a single acceleration factor (8x) was considered, leaving the behavior of the submitted algorithms on higher acceleration factors unknown. Moreover, submitted methods differed not only in architecture, but also in other crucial aspects of the training pipeline such as augmentation and optimization strategies, or ensembling of multiple networks. It is well-known that these factors can preclude any meaningful conclusions from challenge results (Wiesenfarth et al., 2021; Maier-Hein et al., 2018) and the success of a method can potentially be attributed to the wrong component of a system (Eisenmann et al., 2023).
@@ 3.1 Data
Furthermore, we conducted experiments on the publicly available SKM-TEA dataset (Desai et al., 2021b). Similar to the K2S dataset mentioned above, this dataset consists of 155 3-dimensional multi-coil k-space data samples as well as their corresponding segmentations. Using the k-space data and the provided undersampled masks, we selected acceleration factors of 8x and 16x, where 16x corresponds to the maximum acceleration factor for which an undersampling mask is provided by the dataset. The segmentation classes contain the patellar, femoral, and tibial cartilage as well as the meniscus. We use the official data split described in Desai et al. (2021b), which results in a training, test, and validation set of 86, 33, and 36 images, respectively.
@@ 3.2.1 Two-stage methods
CS-Rec. (Donoho, 2006; Lustig et al., 2007): In this approach, first a Compressed Sensing reconstruction (CS-Rec.) is applied to the undersampled image. CS-based reconstruction is a well-established method that is also applied in clinical practice (Donoho, 2006; Lustig et al., 2008; Candès et al., 2006). We use a combination of Total Variation (Block et al., 2007) and L1-Wavelet (Candes et al., 2008) as the optimization objective. In a second step, a segmentation network is trained to segment the anatomical structures from the reconstructed image. We utilize the original code used by Razumov and Dylov (2023) the second-place submission in the K2S challenge (Tolpadi et al., 2023) for the image reconstruction. However, in contrast to the original method, we replace the V-Net segmentation architecture with the same U-Net architecture as NaïveSeg for better comparability to the other methods in our benchmark.
@@ 3.2.2 One-stage methods
TwoDec (Calivá et al., 2020): This method is an adaptation of the TB-Recon method originally proposed by Calivá et al. (2020). It consists of a shared encoder and two decoders that solve the reconstruction and the segmentation, respectively. While the original method employs a V-Net architecture, we opted for a U-Net-like base architecture. We also changed the loss function to include the Dice + Cross-Entropy loss as in all other segmentation models and adjusted the reconstruction function to MSE, as this loss is also utilized for UNet-Rec and RecSeg. This also allows us to isolate the key contribution of (Calivá et al., 2020) – the one encoder, two decoder architecture – and enables a fair comparison to the other methods. Moreover, we observed that a V-Net sometimes led to unstable training on the K2S dataset. As in the original publication we set the weight for the reconstruction loss and the segmentation loss to be equal for the K2S dataset and the SKM-TEA dataset at 8x acceleration, but do an hyperparameter search for the SKM-TEA dataset at 16x acceleration in Section 4.1.
@@ 3.5 Implementation Details
We train and test all methods described in Section 3.2 from scratch on both the K2S and SKM-TEA datasets for all tested acceleration factors using a comparable training pipeline. All segmentation models are trained for 1000 epochs of 250 iterations on an Nvidia Geforce RTX 2080Ti GPU using a batch-size of 2 and the Adam optimizer (Kingma and Ba, 2017). We use a learning rate of for most methods and for RecSeg and TwoDec for training stability. The data-augmentation strategy is identical for all methods and consists of mirroring in the sagittal direction.
@@ 4.1 Segmentation Quality
The segmentation scores for all methods and all acceleration factors for the K2S dataset are shown in Figure 3 (left). The quantitative values for the K2S, and the SKM-TEA datasets are further reported in Tables 1 and 4, respectively. We observed that, overall, the tested methods predicted segmentations of comparable quality and reached similar Dice scores. For both datasets, the difference in Dice score is within one standard deviation between the best and the worst performing method for all tested acceleration factors. The segmentations shown in Figure 2 for the K2S dataset and in Figure 6(b) for the SKM-TEA dataset show examples of the similar segmentation quality between the different methods. It can be seen that the differences in the segmentations are rather small within the same acceleration factor for all tested methods. We further note that the segmentations for all examined methods are of high quality and the errors are often of a similar nature for all methods, like the too large segmentation of the tibia bone as shown in Figure 2.
@@
The method achieving the best Dice scores for the SKM-TEA dataset is Unr-Rec, which achieves the best scores for both tested acceleration factors and is the only method achieving statistically significantly better segmentation results at 16x acceleration. Note that Unr-Rec cannot be tested on the K2S data, because of the 3-dimensional k-space of the MRI images, as explained in Section 3.1.
We found that the performance of methods that combine the reconstruction and segmentation tasks like RecSeg and TwoDec can be dependent on the dataset and the acceleration factor. We observed that both methods are comparable to other methods on the K2S dataset, yet perform worse than other methods on the SKM-TEA dataset for 16x-acceleration (see Table 4). We have analyzed the behavior of the TwoDec method on SKM-TEA with 16x acceleration when using higher weight (higher for the segmentation loss on the validation set
The results are shown in Figure 5. Note that higher segmentation weight leads to higher Dice scores, but the reconstruction scores degrade. For the results in Table 4 we chose , as it provides a good tradeoff between segmentation and reconstruction performance for the SKM-TEA dataset at 16x acceleration. Note that even though we have optimized the segmentation weight hyperparameter for the TwoDec method on the 16x acceleration experiments on SKM-TEA and thereby providing the TwoDec method with an advantage, the results are still worse compared to the other methods.
@@ 4.2 Reconstruction Quality
Tables 2 and 4 provide PSNR and SSIM scores for the reconstruction of different methods. The best reconstructions are created by the two-step UNet-Rec method for the majority of cases. As can be seen in Figures 2 and 6(b), this method is also producing smooth results, which is known to be beneficial for the SSIM and PSNR metrics (Muckley et al., 2021). Lower PSNR scores are reached with the SKM-TEA data in Table 4 compared to the K2S data shown in Table 3. An example image of the reconstructions and the segmentations for the SKM-TEA data with 16x-acceleration factor is shown in Figure 6(b).
@@ 5 Discussion
A disadvantage of the methods that combine the reconstruction with the segmentation task like RecSeg and TwoDec is the dependence on the dataset and acceleration factor. While the methods perform relatively well on the K2S Dataset, their performance degrades for the SKM-TEA dataset at accelerations of 16x as seen in Table 4. The TwoDec method is highly dependent on the selection of the loss weights. Careful tuning of those weights improved the segmentation score, yet the score was still worse than for the other methods.
@@ 6 Conclusions and Future work
In this paper, we have shown that good segmentation quality of undersampled MRI images is feasible even for high acceleration of up to 32x. Contrary to the findings in the K2S challenge (Tolpadi et al., 2023), we have observed that reconstruction methods that enforce data-consistency - either strictly, as in CS-Rec, or implicitly, as in Unr-Rec - lead to the highest segmentation scores. However, we have also shown that the difference between the tested methods can be small, potentially making it difficult to see real-world benefits of the more complex and specialized methods. Experiments have also shown that the performance of some one-stage methods like RecSeg and TwoDec can depend on the dataset and acceleration factor.
@@ Acknowledgements
Funded by the Deutsche Forschungsgemeinschaft (DFG, Ger- man Research Foundation) under Germany’s Excellence Strategy - EXC number 2064/1 - Project number 390727645. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Jan Nikolas Morshuis.

### v2

@@ 2 Related Work
Predicting high-quality segmentations has been the objective in the K2S challenge (Tolpadi et al., 2023), where several methods for segmenting undersampled MRI data have been submitted. While reconstruction fidelity was also analyzed for the methods that did provide reconstructions, it was not used as an evaluation criterion. A drawback of the challenge is that only a single acceleration factor (8) was considered, leaving the behavior of the submitted algorithms on higher acceleration factors unknown. Moreover, submitted methods differed not only in architecture, but also in other crucial aspects of the training pipeline such as augmentation and optimization strategies, or ensembling of multiple networks. It is well-known that these factors can preclude any meaningful conclusions from challenge results (Wiesenfarth et al., 2021; Maier-Hein et al., 2018) and the success of a method can potentially be attributed to the wrong component of a system (Eisenmann et al., 2023).
@@ 3.1 Data
Furthermore, we conducted experiments on the publicly available SKM-TEA dataset (Desai et al., 2021b). Similar to the K2S dataset mentioned above, this dataset consists of 155 3-dimensional multi-coil k-space data samples as well as their corresponding segmentations. Using the k-space data and the provided undersampled masks, we selected acceleration factors of 8 and 16, where 16 corresponds to the maximum acceleration factor for which an undersampling mask is provided by the dataset. The segmentation classes contain the patellar, femoral, and tibial cartilage as well as the meniscus. We use the official data split described in Desai et al. (2021b), which results in a training, test, and validation set of 86, 33, and 36 images, respectively.
@@ 3.2.1 Two-stage methods
CS-Rec. (Donoho, 2006; Lustig et al., 2007): In this approach, first a Compressed Sensing reconstruction (CS-Rec.) is applied to the undersampled image. CS-based reconstruction is a well-established method that is also applied in clinical practice (Donoho, 2006; Lustig et al., 2008; Candès et al., 2006). We use a combination of Total Variation (Block et al., 2007) and L1-Wavelet (Candès et al., 2008) as the optimization objective. In a second step, a segmentation network is trained to segment the anatomical structures from the reconstructed image. We utilize the original code used by Razumov and Dylov (2023) the second-place submission in the K2S challenge (Tolpadi et al., 2023) for the image reconstruction. However, in contrast to the original method, we replace the V-Net segmentation architecture with the same U-Net architecture as NaïveSeg for better comparability to the other methods in our benchmark.
@@ 3.2.2 One-stage methods
TwoDec (Calivá et al., 2020): This method is an adaptation of the TB-Recon method originally proposed by Calivá et al. (2020). It consists of a shared encoder and two decoders that solve the reconstruction and the segmentation, respectively. While the original method employs a V-Net architecture, we opted for a U-Net-like base architecture. We also changed the loss function to include the Dice + Cross-Entropy loss as in all other segmentation models and adjusted the reconstruction function to MSE, as this loss is also utilized for UNet-Rec and RecSeg. This also allows us to isolate the key contribution of (Calivá et al., 2020) – the one encoder, two decoder architecture – and enables a fair comparison to the other methods. Moreover, we observed that a V-Net sometimes led to unstable training on the K2S dataset. As in the original publication we set the weight for the reconstruction loss and the segmentation loss to be equal for the K2S dataset and the SKM-TEA dataset at 8 acceleration, but do an hyperparameter search for the SKM-TEA dataset at 16 acceleration in Section 4.1.
@@ 3.5 Implementation Details
We train and test all methods described in Section 3.2 from scratch on both the K2S and SKM-TEA datasets for all tested acceleration factors using a comparable training pipeline. All segmentation models are trained for 1000 epochs of 250 iterations on an Nvidia Geforce RTX 2080Ti GPU using a batch-size of 2 and the Adam optimizer (Kingma and Ba, 2015). We use a learning rate of for most methods and for RecSeg and TwoDec for training stability. The data-augmentation strategy is identical for all methods and consists of mirroring in the sagittal direction.
@@ 4.1 Segmentation Quality
The segmentation scores for all methods and all acceleration factors for the K2S dataset are shown in Figure 3 (left). The quantitative values for the K2S, and the SKM-TEA datasets are further reported in Tables 1 and 4, respectively. We observed that, overall, the tested methods predicted segmentations of comparable quality and reached similar Dice scores. For both datasets, the difference in Dice score is within one standard deviation between the best and the worst performing method for all tested acceleration factors. The segmentations shown in Figure 2 for the K2S dataset and in Figure 6 for the SKM-TEA dataset show examples of the similar segmentation quality between the different methods. It can be seen that the differences in the segmentations are rather small within the same acceleration factor for all tested methods. We further note that the segmentations for all examined methods are of high quality and the errors are often of a similar nature for all methods, like the too large segmentation of the tibia bone as shown in Figure 2.
@@
The method achieving the best Dice scores for the SKM-TEA dataset is Unr-Rec, which achieves the best scores for both tested acceleration factors and is the only method achieving statistically significantly better segmentation results at 16 acceleration. Note that Unr-Rec cannot be tested on the K2S data, because of the 3-dimensional k-space of the MRI images, as explained in Section 3.1.
We found that the performance of methods that combine the reconstruction and segmentation tasks like RecSeg and TwoDec can be dependent on the dataset and the acceleration factor. We observed that both methods are comparable to other methods on the K2S dataset, yet perform worse than other methods on the SKM-TEA dataset for 16 acceleration (see Table 4). We have analyzed the behavior of the TwoDec method on SKM-TEA with 16 acceleration when using higher weight (higher for the segmentation loss on the validation set
The results are shown in Figure 5. Note that higher segmentation weight leads to higher Dice scores, but the reconstruction scores degrade. For the results in Table 4 we chose , as it provides a good tradeoff between segmentation and reconstruction performance for the SKM-TEA dataset at 16 acceleration. Note that even though we have optimized the segmentation weight hyperparameter for the TwoDec method on the 16 acceleration experiments on SKM-TEA and thereby providing the TwoDec method with an advantage, the results are still worse compared to the other methods.
@@ 4.2 Reconstruction Quality
Tables 2 and 4 provide PSNR and SSIM scores for the reconstruction of different methods. The best reconstructions are created by the two-step UNet-Rec method for the majority of cases. As can be seen in Figures 2 and 6, this method is also producing smooth results, which is known to be beneficial for the SSIM and PSNR metrics (Muckley et al., 2021). Lower PSNR scores are reached with the SKM-TEA data in Table 4 compared to the K2S data shown in Table 3. An example image of the reconstructions and the segmentations for the SKM-TEA data with 16-acceleration factor is shown in Figure 6.
@@ 5 Discussion
A disadvantage of the methods that combine the reconstruction with the segmentation task like RecSeg and TwoDec is the dependence on the dataset and acceleration factor. While the methods perform relatively well on the K2S Dataset, their performance degrades for the SKM-TEA dataset at accelerations of 16 as seen in Table 4. The TwoDec method is highly dependent on the selection of the loss weights. Careful tuning of those weights improved the segmentation score, yet the score was still worse than for the other methods.
@@ 6 Conclusions and Future work
In this paper, we have shown that good segmentation quality of undersampled MRI images is feasible even for high acceleration of up to 32. Contrary to the findings in the K2S challenge (Tolpadi et al., 2023), we have observed that reconstruction methods that enforce data-consistency - either strictly, as in CS-Rec, or implicitly, as in Unr-Rec - lead to the highest segmentation scores. However, we have also shown that the difference between the tested methods can be small, potentially making it difficult to see real-world benefits of the more complex and specialized methods. Experiments have also shown that the performance of some one-stage methods like RecSeg and TwoDec can depend on the dataset and acceleration factor.
@@ Acknowledgements
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC number 2064/1 - Project number 390727645. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Jan Nikolas Morshuis.